Questions 33-37 An agriculiural ecooit wants to study the relatioship betwen the aoun o Sertiliner snd and the yield. In Y bo the yield measured in kilograms per acre, and X be thr maat of แ¡tugm fortika measured in kilograms per arre, She collects a random sample o plots and estimates a ssion model of OUTPUT Multiple R .0887 R Square 2379 1518550052 2380166120491 287.29 18.43 15.59 51.15 23.43 xVariable 1 33 Find the significance levels at which the coefticient of ftlis statisticaly sigaificnt (a) 1% (b) 10% (e) 1%,5% and 10% 34 The cmist is inteeestel in testing yotes Ho: ersH&.Find the eof this bypothesis testing (๑) 0.000% (b) 2.978% (e) 4.133% (d) 840 )17361 a5 If a fokg of e, what i the espected amount of yield er ace aeding to the gn s () 773.3 (b) 758.29 () 605.21 (d) 894.29 (e) SOLOS Q36 If a farme does sot use any Setiliers, what is the esxpercted anmd of yel er ace anc eding bo the regrsion wults
Answer #1

33)

option e is correct

34)

test statisic =(4.71-5)/0.31=-0.94

for abvoe test statistic p value =17.361%

option E

35)

expected yield=287.29+4.71*100=758.29

option B

36)

expected amount =287.29